OpenStudy (anonymous):

Please solve Ill give medal! -9 = 5 over 17n

OpenStudy (anonymous):

you wanna solve for n multiply 17n both sides.

OpenStudy (anonymous):

-5 over 153? @mathessentials

OpenStudy (anonymous):

\[-9=\frac{ 5 }{ 17n }\]multiply both sides by 17n

OpenStudy (anonymous):

-17 over 45 ?

OpenStudy (anonymous):

\[-9 \ \times 17n=5\]

OpenStudy (anonymous):

because 17n/17n=1 and cancels

OpenStudy (anonymous):

well i have two choices with 5 which are -153/5 and -5/153

OpenStudy (anonymous):

I see

OpenStudy (anonymous):

if you take this and divide from both sides all that's not n\[-9 \ \times 17n=5\]you get\[n=\frac{5}{-9 \ \times 17}\]